Th minimum possible electric force betwn 2 charges kept 3 m apart?

Grade:12

1 Answers

Shubhangi Jha
6 years ago
Take charge as 1.6×10^-19C and solve using fomula F=kq^2/r^2F=9×10^9*(1.6×10^-19)^2/3*3 =2.56×10^-29N